a. Analyse the cross price elasticity of demand for wheat and rice when a change in the price of wheat from Rs.70 to Rs. 90 results in a change in the quantity demand for wheat from kg 3000 to kg 5000 in the market. Interpret the value of the coefficient.


1
Expert's answer
2021-04-21T13:22:08-0400

"E_{wr} = \frac{\frac{\Delta Q_w}{Q_w}}{\frac{\Delta P_r}{P_r}} = \frac{\frac{2000}{3000}}{\frac{20}{70}} = \frac{2/3}{2/7} = \frac{7}{3} = 2,33"

Since the cross-price elasticity is positive those goods are substitutes which means that the demand for one good increases when the price for the substitute good increases